OpenStudy (anonymous):

find Cos 2A, cos A= √3/4

OpenStudy (kropot72):

Use the following formula: \[\cos 2A=2\cos ^{2}A-1\] \[\cos 2A=(2\times (\frac{\sqrt{3}}{4})^{2})-1=you\ can\ calculate\]

OpenStudy (anonymous):

cos2A=2cos2A−1 put cos A= √3/4 so cos2A=2*(√3/4)*(√3/4)-1